Thursday, September 7. Bruce noticed that the shots per scan we getting noisy (not stable around 170) and slowly decreasing. The system was restarted at approximately 0810 PDT with a 5 minute warmup and the BSU was reinitialized.

This is the second time that this has occurred - both times the lidar had been operating for multiple days. The first event happened on Sept 2.

LabVIEW restart
Clear and sunny today.
![]()

Some decent gust structures this afternoon in the REAL images. Here is one example:
![]()
![]()
![]()
![]()
![]()
and some corresponding wind retrievals:
![]()
![]()
![]()
![]()
Here's another sequence:
![]()
![]()
![]()
![]()
![]()
![]()
![]()
and some corresponding wind fields:
![]()
![]()
![]()
![]()
But, sometimes conditions become "quiet" (meaning, no gust structures):
![]()
![]()
Shane visited site from about 8:15 PM until about 10:00 PM PDT. Checked mirror downstream of Raman cell. Inclusion does not look any more advanced then prior photos.

0 Z on 8 Sept 500 mb chart:
